OpenStudy (anonymous):

How do I simplify this? (4y/(y+2))+(8/(y+2)) Thanks in advance!

OpenStudy (anonymous):

I got 4x+8/(y+2) and it said it was wrong. I am really confused.

hero (hero):

I see. Hang on. Let me re-post one that is even more demonstrative.

hero (hero):

In other words, you need to factor something out of the numerator to so that it is in simplest form.

hero (hero):

Oh but, now when you simplify it, you'll have to do something similar to this:

hero (hero):

\[\frac{8x }{x + 2} + \frac{16}{x + 2} = \frac{8x + 16}{x + 2} = \frac{8(x + 2)}{(x +2)} = 8\]

hero (hero):

What you have to do is similar to that above.

hero (hero):

You have to cancel something from the top and bottom.

OpenStudy (anonymous):

So it would equal 4?

hero (hero):

Yes, correct....finally

OpenStudy (anonymous):

That took me longer than it should have. Thank you!

hero (hero):

You must familiarize yourself more with it.